# ADPO: Anchored Direct Preference Optimization A Unified Framework from Pairwise to Listwise Preferences

### v5

@@ 3.1 Core Formulation: Anchored Distribution as the Protagonist
    DPO (binary preferences). For $|S_{x}|=2$ with Bradley–Terry targets $q(i\mid\{i,j\})=\sigma(\beta_{r}\Delta R)$, Eq. (5) reduces to DPO [1] when $\beta=\beta_{r}$ and $\pi_{\mathrm{ref}}$ is uniform.
Remark 3.3 (Connection to AWR/AWAC/MPO).
ADPO also connects to offline RL methods at the $\tilde{p}-$distribution level; see Appendix A for details on this partial equivalence.
@@
Table 2: ADPO as a unifying projection framework. Different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau_{\mathrm{anc}}$ recover major learning paradigms. Key: Optimization matches the anchored distribution $\tilde{p}_{\theta}=\mathrm{softmax}((s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}})$ to the target $q$; the corresponding policy $\pi_{\theta}$ satisfies $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$ when the optimum is reachable (Prop. 3.4).
@@
KD (standard) $\mathrm{softmax}(z_{T}/T)$ uniform $T$ $T^{2}\,\mathrm{CE}(q_{T},\mathrm{softmax}(s/T))$
@@
Proposition 3.4 (Closed-form optimal solution).
Suppose $S_{x}$ covers the full action/class space and $q(\cdot\mid S_{x})\succ 0$. (Discrete case) If the policy $\pi_{\theta}$ has sufficient expressivity (i.e., the parametrization can represent arbitrary log-odds over $S_{x}$), then the distribution-level optimum of $\operatorname*{arg\,min}_{\theta}\mathrm{KL}(q\|\tilde{p}_{\theta})$ satisfies
$\pi_{\theta}(i\mid x)\;\propto\;\pi_{\mathrm{ref}}(i\mid x)\cdot q(i\mid x)^{\tau_{\mathrm{anc}}}.$ (8)
(Continuous case) For continuous action spaces, assume $q$, $\pi_{\theta}$, and $\pi_{\mathrm{ref}}$ admit densities with respect to a common base measure (e.g., Lebesgue measure). Define the anchored distribution via the density ratio:
$\tilde{p}_{\theta}(a)\;\propto\;\exp\!\Big(\frac{\log\pi_{\theta}(a)-\log\pi_{\mathrm{ref}}(a)}{\tau_{\mathrm{anc}}}\Big)\;=\;\Big(\frac{\pi_{\theta}(a)}{\pi_{\mathrm{ref}}(a)}\Big)^{1/\tau_{\mathrm{anc}}}.$
Setting $\tilde{p}_{\theta}=q$ yields the closed-form solution
$\pi_{\theta}(a)\;\propto\;\pi_{\mathrm{ref}}(a)\cdot q(a)^{\tau_{\mathrm{anc}}},$
which requires: (i) overlapping support: $q(a)>0\Rightarrow\pi_{\mathrm{ref}}(a)>0$ almost everywhere; (ii) integrability: $\int\pi_{\mathrm{ref}}(a)\cdot q(a)^{\tau_{\mathrm{anc}}}\,da<\infty$ to ensure normalizability. Under these conditions, the solution is unique up to the normalization constant (equivalently, the Radon–Nikodym derivative $d\pi_{\theta}/d\mu$ is determined by the multiplicative form, where $\mu$ is the base measure).
Proof (discrete): $\tilde{p}_{\theta}=\mathrm{softmax}\big((s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}\big)=q$ if and only if $(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}=\log q+c\mathbf{1}$ for some constant $c$ (reflecting the shift-invariance of softmax), which rearranges to $\pi_{\theta}\propto\pi_{\mathrm{ref}}\,q^{\tau_{\mathrm{anc}}}$. The constant $c$ is absorbed into the proportionality constant. ∎
Implication and reachability. Equation (8) strictly characterizes the "anchored projection" as a multiplicative fusion of reference and target when the optimum is reachable. In discrete finite-action settings with tabular or sufficiently expressive function approximation, this optimum is exactly attainable. However, in structured settings (e.g., autoregressive language models with shared parameters across tokens, or policy networks with limited capacity), the optimum may be unreachable, and gradient descent performs an approximate projection in the parametric family. The KD, MaxEnt, and AWR special cases in Proposition 3.2 all follow directly from (8).
Listwise subset constraints and convexity. When $S_{x}$ is only a subset (pairwise/listwise ranking), (8) holds within each group, but the global $\pi_{\theta}$ must simultaneously satisfy multiple group constraints. Crucially, $\mathcal{L}_{\mathrm{ADPO}}$ is convex with respect to the anchored logits $u=(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ within each group (as shown in Theorem 3.8), but the optimization over shared parameters $\theta$ is typically non-convex. The group-wise closed-form solution provides a target, but there may not exist a parameter $\theta$ that simultaneously achieves $\tilde{p}_{\theta}(\cdot\mid S_{x})=q(\cdot\mid S_{x})$ for all groups (except in the tabular case or with unlimited expressivity). Thus, in practice, ADPO performs a parameterized fitting that minimizes the expected group-wise KL, balancing the constraints across all training groups.
By centering the coordinate system at $\pi_{\mathrm{ref}}$ before applying softmax, ADPO provides groupwise shift invariance (Lemma 2.1) and Fisher-metric regularization (Lemma 3.5)—properties absent in standard KD or unanchored DPO. Table 2 summarizes the unification.
Mirror descent = KL proximal step.
If one directly optimizes over the anchored logits $u$ (e.g., tabular setting), each gradient step is equivalent to a KL-trust-region proximal problem (Theorem 3.7):
$u_{t+1}=\arg\min_{u}\ \langle\nabla\mathcal{L}(u_{t}),\,u\rangle+\tfrac{1}{\eta}\mathrm{KL}(\tilde{p}(u_{t})\|\tilde{p}(u)),$
whose solution is the multiplicative-weights update $\tilde{p}_{t+1}\propto\tilde{p}_{t}\odot\exp(-\eta(\tilde{p}_{t}-q))$. This exhibits ADPO as a Bregman-proximal method in distribution space.
Dynamic anchoring: monotonicity within fixed-anchor phases.
When the anchor $\pi_{\mathrm{ref}}$ is held fixed, the loss $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ (where $A(u)=\log\sum_{j}e^{u_{j}}$) is strictly convex on the subspace $\mathbf{1}^{\perp}$. Under either (i) local analysis near $u^{\star}$ or (ii) bounded-domain constraints $\tilde{p}_{\theta}\in[\varepsilon,1-\varepsilon]^{K}$, $\mathcal{L}$ is $\mu$-strongly convex on $\mathbf{1}^{\perp}$, guaranteeing linear convergence $\mathcal{L}(u_{t+1})-\mathcal{L}(u^{\star})\leq(1-\mu\eta)(\mathcal{L}(u_{t})-\mathcal{L}(u^{\star}))$ within each fixed-anchor phase (Theorem 3.8). When the anchor updates ($\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$), the coordinate system changes, and we do not claim monotonicity across anchor updates—only within each phase.
Standing assumptions.
Throughout this work, we maintain the following conditions unless otherwise noted:
  1. (i)
    Support overlap (discrete): For discrete action/candidate spaces, we assume $q(i\mid x)>0\Rightarrow\pi_{\mathrm{ref}}(i\mid x)>0$ for all $i\in S_{x}$, ensuring that anchored logits $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ are well-defined.
  2. (ii)
    Absolute continuity (continuous): For continuous action spaces, we assume $q$, $\pi_{\theta}$, and $\pi_{\mathrm{ref}}$ admit densities with respect to a common base measure, with overlapping support in the same sense as (i).
  3. (iii)
    Local analysis vs. global bounds: Fisher-metric trust region results (Lemma 3.5, Corollary 3.6) hold locally around $u^{\star}$ where $\tilde{p}_{\theta}(u^{\star})=q$. For global strong convexity, we either work in a neighborhood of $u^{\star}$ or assume the optimization is constrained to a bounded domain where $\tilde{p}_{\theta}(u)_{i}\in[\varepsilon,1-\varepsilon]$ (Theorem 3.8).
  4. (iv)
    Sequence models: For autoregressive sequence generation, we use length-normalized (per-token average) log-probabilities $s$ and $s^{\mathrm{ref}}$ as group-level scores, and apply the softmax Fisher geometry at the sequence-level distribution over candidates.
  5. (v)
    Expressivity: Closed-form optimal solutions (Proposition 3.4) assume the policy parametrization $\pi_{\theta}$ has sufficient expressivity to represent the target multiplicative form $\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$. In practice (e.g., shared-parameter networks, autoregressive models), we perform an approximate projection in the parametric family.
3.2 Pairwise Instantiation
Notation for this subsection: $\beta$ = student inverse temperature, $\beta_{r}$ = teacher/reference inverse temperature, $q_{ij}$ = soft preference probability that $i\succ j$.
For a candidate pair $(i,j)$ with soft teacher preference $q_{ij}\in(0,1)$ that $i$ is preferred, the ADPO pairwise loss is:
$\ell_{ij}^{\text{Soft-DPO}}=\log\left(1+e^{\beta(\Delta_{\theta}-\Delta_{\text{ref}})}\right)-q_{ij}\,\beta(\Delta_{\theta}-\Delta_{\text{ref}}),$ (9)
where $\beta>0$ is student temperature, $\Delta_{\theta}=s_{i}-s_{j}$, and $\Delta_{\text{ref}}=s_{i}^{\text{ref}}-s_{j}^{\text{ref}}$ with $s_{i}=\log\pi_{\theta}(y_{i}|x)$.
Key properties:
  • Bayes-optimal matching: Gradient vanishes when $\sigma(\beta(\Delta_{\theta}-\Delta_{\text{ref}}))=q_{ij}$.
  • Soft weighting: Uncertain pairs ($q_{ij}\approx 0.5$) contribute less gradient.
  • Reference anchoring: Relative updates $\Delta_{\theta}-\Delta_{\text{ref}}$ are invariant to groupwise shifts.
Temperature alignment: The pairwise formulation uses inverse temperature $\beta$, which relates to the listwise anchoring temperature via $\beta=1/\tau_{\mathrm{anc}}$. Specifically, $\tilde{p}_{\theta}(i)=\sigma\!\big((\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau_{\mathrm{anc}}\big)=\sigma\!\big(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}})\big)$.
3.3 Listwise Instantiation
Notation for this subsection: $\tau_{\mathrm{anc}}$ = anchor temperature, $\beta_{r}$ = teacher inverse temperature for forming $q$ from rewards, $\tilde{R}_{i}$ = observed rewards, $\hat{R}_{i}$ = transformed rewards.
For group $S_{x}=\{y_{1},\ldots,y_{P}\}$, the ADPO listwise loss is:
$\mathcal{L}_{\text{group}}^{\text{ref}}=\mathbb{E}_{x,S_{x}}\left[-\sum_{i\in S_{x}}q(i\mid S_{x})\log\tilde{p}_{\theta}(i\mid S_{x})\right],$ (10)
where
$\tilde{p}_{\theta}(i\mid S_{x})=\frac{\exp\left((s_{i}-s_{i}^{\text{ref}})/\tau_{\mathrm{anc}}\right)}{\sum_{j\in S_{x}}\exp\left((s_{j}-s_{j}^{\text{ref}})/\tau_{\mathrm{anc}}\right)},$ (11)
and teacher target $q(i\mid S_{x})\propto\exp(\hat{R}_{i}/\beta_{r})$ uses transformed rewards $\hat{R}_{i}$ via: (a) raw: $\hat{R}_{i}=\tilde{R}_{i}$; (b) rank-based Gaussian transform; or (c) KDE-CDF-Logit: $\hat{R}_{i}=\mathrm{logit}(\hat{F}(\tilde{R}_{i}))$ where $\hat{F}$ is the KDE-estimated CDF.
Equivalent expansion. The anchored listwise cross-entropy (10) can be equivalently written as:
$\sum_{i\in S_{x}}q(i\mid S_{x})\log\tilde{p}_{\theta}(i\mid S_{x})=-\frac{1}{\tau_{\mathrm{anc}}}\sum_{i\in S_{x}}q(i\mid S_{x})\big(s_{i}-s_{i}^{\mathrm{ref}}\big)+\log\!\sum_{j\in S_{x}}\exp\!\left(\frac{s_{j}-s_{j}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}\right).$ (12)
This decomposition makes explicit that the anchored listwise cross-entropy consists of a linear matching term $\langle q,(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}\rangle$ and a log-sum-exp normalization. It directly reveals the connection to the convex conjugate form of the softmax, underlying the implicit trust-region regularization discussed below. Since $\mathrm{KL}(q\|\tilde{p}_{\theta})=\mathcal{L}_{\text{group}}^{\text{ref}}+H(q)$ and $H(q)$ is parameter-independent, minimizing (12) is equivalent to minimizing the KL divergence from the teacher distribution $q$ to the student’s anchored distribution $\tilde{p}_{\theta}$.
3.4 Formal Statements of Core Results
We now provide formal statements and proofs for the results sketched in Section 3.
@@ 3.4 Formal Statements of Core Results
Let $u=(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ and $\tilde{p}(u)=\mathrm{softmax}(u)$. Let $u^{\star}=\log q+c\mathbf{1}$ so that $\tilde{p}(u^{\star})=q$. Define $\bar{\delta}=(u-u^{\star})-\langle q,u-u^{\star}\rangle\mathbf{1}$, the $q$-centered deviation. Then, as $u\to u^{\star}$,
$\mathrm{KL}\!\big(q\,\|\,\tilde{p}(u)\big)\;=\;\tfrac{1}{2}\,\bar{\delta}^{\top}\!\big(\operatorname{Diag}(q)-qq^{\top}\big)\,\bar{\delta}\;+\;o(\|\bar{\delta}\|^{2}).$
The curvature is given by the softmax Fisher metric $\operatorname{Diag}(q)-qq^{\top}$, which is independent of the anchor choice and has a null-space along $\mathbf{1}$.
@@
Let $A(u)=\log\sum_{j}e^{u_{j}}$ so that $\mathcal{L}(u)=A(u)-\langle q,u\rangle$. Then $\nabla A(u)=\tilde{p}_{\theta}(\cdot;u)$ and $\nabla^{2}A(u)=\operatorname{Diag}(\tilde{p}_{\theta})-\tilde{p}_{\theta}\tilde{p}_{\theta}^{\top}$. A second-order Taylor expansion of $\mathcal{L}$ at $u^{\star}$ with $\tilde{p}_{\theta}(\cdot;u^{\star})=q$ yields the stated quadratic form. The centering removes the null-space along $\mathbf{1}$, reflecting the invariance of softmax to additive shifts. ∎
Corollary 3.6 (Spectral bounds).
Let $\mu(q)>0$ denote the smallest nonzero eigenvalue of $\operatorname{Diag}(q)-qq^{\top}$. In the quotient space $\mathbb{R}^{K}/\mathrm{span}\{\mathbf{1}\}$, the objective is strongly convex. Locally around $u^{\star}$,
$\mathrm{KL}(q\|\tilde{p}(u))\;\geq\;\frac{\mu(q)}{2}\,\|\bar{\delta}\|^{2},$
where $\bar{\delta}=(u-u^{\star})-\langle q,u-u^{\star}\rangle\mathbf{1}$. By Pinsker’s inequality, $\|\tilde{p}(u)-q\|_{1}\leq\sqrt{2\mathrm{KL}(q\|\tilde{p}(u))}$.
For $K=2$: $\mu(q)=2q(1-q)$. For $K>2$, when $\min_{i}q_{i}\geq\varepsilon>0$, there exists a constant $c(\varepsilon,K)>0$ such that $\mu(q)\geq c(\varepsilon,K)$ on the quotient space. In particular, for any $v$ satisfying $\langle q,v\rangle=0$,
$\operatorname{Var}_{q}[v]=\sum_{i}q_{i}v_{i}^{2}\geq\varepsilon\|v\|^{2}.$
Geometric interpretation.

### v6

@@ 3.2 Unified ADPO Objective and Its Special Cases
For $|S_{x}|{=}2$, let $\Delta_{\theta}=\ell_{i}-\ell_{j}$ and $\Delta_{\mathrm{ref}}=\ell_{i}^{\mathrm{ref}}-\ell_{j}^{\mathrm{ref}}$. With a Bradley–Terry target $q_{ij}\in(0,1)$, (4) reduces to:
$\boxed{\mathcal{L}_{ij}^{\text{ADPO}}=\log\!\big(1+\exp[(\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau_{\mathrm{anc}}]\big)-q_{ij}\,(\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau_{\mathrm{anc}},}$ (7)
which recovers the standard DPO loss [1] when $\Delta_{\mathrm{ref}}=0$ (uniform anchor) and $q_{ij}\in\{0,1\}$ (hard labels). When $q_{ij}\in(0,1)$ (soft labels from noisy annotations), this becomes "soft DPO" that naturally handles label uncertainty. When $\Delta_{\mathrm{ref}}\neq 0$ (non-uniform anchor), we obtain "anchored DPO" with improved stability.
Gradient.
The gradient of the pairwise loss is:
$\frac{\partial\mathcal{L}_{ij}^{\mathrm{ADPO}}}{\partial\theta}=\frac{1}{\tau_{\mathrm{anc}}}\big(\sigma[(\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau_{\mathrm{anc}}]-q_{ij}\big)(\nabla_{\theta}\ell_{i}-\nabla_{\theta}\ell_{j}),$
where $\sigma(\cdot)$ is the sigmoid function.
Listwise (general) case.
For candidate sets $S_{x}=\{y_{1},\dots,y_{P}\}$ with $P>2$, the ADPO listwise objective is:
$\boxed{\mathcal{L}_{\text{group}}^{\text{ADPO}}=-\sum_{i\in S_{x}}q(i|S_{x})\log\tilde{p}_{\theta}(i|S_{x}),\qquad\tilde{p}_{\theta}(\cdot|S_{x})=\mathrm{softmax}(\bar{u}),}$ (8)
with gradient
$\frac{\partial\mathcal{L}_{\mathrm{group}}^{\text{ADPO}}}{\partial\ell_{i}}=\frac{1}{\tau_{\mathrm{anc}}}\big(\tilde{p}_{\theta}(i|S_{x})-q(i|S_{x})\big).$ (9)
Setting P=2P{=}2 recovers the binary case (7), showing that ADPO continuously extends DPO from pairwise to listwise soft supervision.
$q$-centering: A principled use of shift-invariance.
Recall from Section 3.1 that softmax is shift-invariant: $\mathrm{softmax}(u)=\mathrm{softmax}(u+c\mathbf{1})$ for any constant $c$. We exploit this property by choosing $c=-\sum_{j}q_{j}u_{j}$ (the negative $q$-weighted mean), yielding:
$\boxed{\bar{u}_{i}=u_{i}-\sum_{j\in S_{x}}q(j|S_{x})\,u_{j},\qquad u_{i}=\frac{\ell_{i}-\ell_{i}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}.}$ (10)
This $q$-centering is always applied in our algorithm (Algorithm 1, both modes) and serves three purposes: numerical stability (centers logits near zero); geometric consistency (removes null-space along $\mathbf{1}$ in Fisher metric, Section 3.3); theoretical alignment (consistent with $q$-centered deviation $\bar{\delta}$ in Lemma 3.6).
Our final anchored distribution is thus:
$\tilde{p}_{\theta}(i|S_{x})=\mathrm{softmax}(\bar{u})_{i}=\frac{\exp(\bar{u}_{i})}{\sum_{j\in S_{x}}\exp(\bar{u}_{j})}.$ (11)
Remark 3.3 (Optional: Fisher variance scaling).
For extremely large $P$ (e.g., $P\geq 16$), one can additionally apply Fisher normalization:
$\widehat{u}_{i}=\frac{\bar{u}_{i}}{\sqrt{\mathrm{Var}_{q}[u]+\varepsilon}},\qquad\mathrm{Var}_{q}[u]=\sum_{j}q_{j}u_{j}^{2}-\Big(\sum_{j}q_{j}u_{j}\Big)^{2}.$ (12)
This "whitens" the gradient but changes the effective temperature. We do not use this in our experiments and recommend it only as a troubleshooting step for very large $P$.
Unifying major learning paradigms.
Different choices of $(q,\pi_{\mathrm{ref}},\tau_{\mathrm{anc}})$ recover major learning paradigms:
Proposition 3.4 (ADPO unifies common paradigms).
Specific choices of $(q,\pi_{\mathrm{ref}},\tau_{\mathrm{anc}})$ recover:
  • Supervised Fine-Tuning (SFT): $q$ from labeled data, $\pi_{\mathrm{ref}}$ uniform, $\tau_{\mathrm{anc}}=1$.
  • Knowledge Distillation (KD): $q=\mathrm{softmax}(z_{T}/T)$ from teacher logits $z_{T}$, $\pi_{\mathrm{ref}}$ uniform, $\tau_{\mathrm{anc}}=T$.
  • Maximum-Entropy RL: $q(a|s)\propto\exp(Q/\alpha)$ from Q-values, $\pi_{\mathrm{ref}}$ uniform, $\tau_{\mathrm{anc}}=1$.
  • Direct Preference Optimization (DPO): $|S_{x}|=2$, $q(i)=\sigma(\beta_{r}\Delta R)$ from reward differences, $\pi_{\mathrm{ref}}$ uniform.
Verification.
Each case follows from Proposition 3.2 by substituting the specified $(q,\pi_{\mathrm{ref}},\tau_{\mathrm{anc}})$. ∎
Table 1 summarizes this unification.
@@
Table 1: ADPO as a unifying projection framework. Different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau_{\mathrm{anc}}$ recover major learning paradigms. Key: Optimization matches the anchored distribution $\tilde{p}_{\theta}=\mathrm{softmax}((\ell-\ell^{\mathrm{ref}})/\tau_{\mathrm{anc}})$ to the target $q$; the corresponding policy $\pi_{\theta}$ satisfies $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$ when the optimum is reachable (Prop. 3.2). Caution: When anchor and target both equal the teacher ($\pi_{\mathrm{ref}}=\pi_{T},q=\pi_{T}$), the optimum becomes $\pi_{\theta}\propto\pi_{T}^{1+\tau_{\mathrm{anc}}}$, which sharpens the teacher distribution and may degrade calibration; use small $\tau_{\mathrm{anc}}$ or add auxiliary regularization if needed.
@@
KD (standard) $\mathrm{softmax}(z_{T}/T)$ uniform $T$ $T^{2}\,\mathrm{CE}(q_{T},\mathrm{softmax}(\ell/T))$
@@
Warning: Sharpens teacher; see caption.
3.3 Fisher Geometry: An Implicit Trust Region
The anchored softmax structure induces a natural Riemannian geometry on the probability simplex. Let $\tilde{p}_{\theta}$ (hereafter $p$ for brevity) denote the current anchored distribution. The softmax Fisher information matrix is:
$F\;=\;\operatorname{Diag}(p)-pp^{\top}\;\in\;\mathbb{R}^{P\times P},\qquad F\mathbf{1}{=}0,$ (13)
where $F\mathbf{1}=0$ reflects the normalization constraint and shift-invariance. Near the optimum $u^{\star}$ (where $\tilde{p}_{\theta}(u^{\star})=q$), the KL divergence admits a second-order expansion:
$\mathrm{KL}(q\|\tilde{p}(u))\;\approx\;\tfrac{1}{2\tau_{\mathrm{anc}}^{2}}\,\bar{\delta}^{\top}(\operatorname{Diag}(q)-qq^{\top})\bar{\delta},\qquad\bar{\delta}=(u-u^{\star})-\langle q,u-u^{\star}\rangle\mathbf{1},$ (14)
where the $1/\tau_{\mathrm{anc}}^{2}$ scaling governs the local curvature (see Figure 1). This creates an implicit trust region centered at the target $q$, independent of the anchor $\pi_{\mathrm{ref}}$—the anchor defines the coordinate system, while the Fisher metric determines the geometry.
Figure 1: Fisher geometry as implicit trust region. The Fisher metric $\operatorname{Diag}(q)-qq^{\top}$ defines a quadratic bowl in anchored coordinates $u=(\ell-\ell^{\mathrm{ref}})/\tau_{\mathrm{anc}}$, centered at target $q$. Smaller $\tau_{\mathrm{anc}}$ yields tighter curvature ($\propto 1/\tau_{\mathrm{anc}}^{2}$).
Natural gradient and pullback metric.
The natural gradient $\delta u_{\text{nat}}=F^{+}(q-p)$ (where $F^{+}$ is the Moore–Penrose pseudoinverse) gives the steepest descent direction in the Fisher metric [27]. In parameter space, the pullback metric is:
$\mathsf{G}(\theta;S_{x})=J\,F\,J^{\top}=\sum_{i}p_{i}g_{i}g_{i}^{\top}-\Big(\sum_{i}p_{i}g_{i}\Big)\Big(\sum_{i}p_{i}g_{i}\Big)^{\top},\quad J=[g_{1},\ldots,g_{P}],\ g_{i}=\nabla_{\theta}\ell_{i}.$ (15)
This is exactly the Fisher information matrix used in TRPO [23], but ADPO induces it implicitly through the anchored softmax—no explicit second-order optimization is required.
Remark 3.5 (Theory vs practice).
Our implementation (Algorithm 1, both modes) uses standard first-order optimization (Adam) on the ADPO loss (4). The Fisher geometry provides theoretical insight into why this works well, not how we implement it.
3.4 Mirror-Descent Equivalence and Convergence
The ADPO update can be viewed as mirror descent with KL Bregman divergence. Within each group, the loss $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ (where $A(u)=\log\sum_{j}e^{u_{j}}$) is convex in anchored logits $u=(\ell-\ell^{\mathrm{ref}})/\tau_{\mathrm{anc}}$.
@@ 3.4 Mirror-Descent Equivalence and Convergence
Near the optimum $u^{\star}$ (where $\tilde{p}_{\theta}(u^{\star})=q$), writing $\bar{\delta}=(u-u^{\star})-\langle q,u-u^{\star}\rangle\mathbf{1}$,
$\mathrm{KL}\big(q\|\tilde{p}(u)\big)\;=\;\tfrac{1}{2\tau_{\mathrm{anc}}^{2}}\,\bar{\delta}^{\top}\big(\operatorname{Diag}(q)-qq^{\top}\big)\bar{\delta}\;+\;o(\|\bar{\delta}\|^{2}).$ (16)
@@
Second-order Taylor expansion of $\mathcal{L}$ at $u^{\star}$ with $\nabla^{2}A(u)=\operatorname{Diag}(\tilde{p})-\tilde{p}\,\tilde{p}^{\top}$ yields the stated form. The $q$-centering removes the null-space along $\mathbf{1}$. ∎
Theorem 3.7 (Mirror-descent equivalence).
The ADPO update with step size $\eta$ satisfies:
$u_{t+1}=\arg\min_{u}\big\langle\nabla\mathcal{L}(u_{t}),\,u\big\rangle+\frac{1}{\eta}\,\mathrm{KL}\!\big(\tilde{p}(u_{t})\,\|\,\tilde{p}(u)\big),$
inducing the multiplicative rule $\tilde{p}_{t+1}\propto\tilde{p}_{t}\odot\exp(-\eta(\tilde{p}_{t}-q))$ on distributions. See Appendix for full derivation.
Corollary 3.8 (Dynamic anchor as proximal trust region).
When $\pi_{\mathrm{ref}}=\pi_{\text{old}}$ (dynamic anchor), each step controls $\mathrm{KL}(\tilde{p}_{\text{old}}\|\tilde{p}_{\text{new}})$ in anchored coordinates, mirroring TRPO’s $\mathrm{KL}(\pi_{\text{old}}\|\pi_{\text{new}})$ constraint with groupwise shift invariance. The temperature $\tau_{\mathrm{anc}}$ scales curvature as $1/\tau_{\mathrm{anc}}^{2}$, acting as a radius knob.
3.5 Anchor Strategies: Offline vs Online
The flexibility of ADPO’s anchoring mechanism enables different instantiations depending on the choice of anchor policy $\pi_{\mathrm{ref}}$. Because the Fisher metric depends only on the target $q$ (not on $\pi_{\mathrm{ref}}$), we can cleanly separate the choice of anchor from the learning objective, leading to two main paradigms.
Offline ADPO: Fixed or no anchor.
In offline settings (SFT, KD, preference learning from fixed datasets), the anchor is either:
  • No anchor (uniform): Set $\pi_{\mathrm{ref}}$ uniform, equivalent to $\ell^{\mathrm{ref}}_{i}=0$. This recovers:
    • SFT: $q$ from labeled data, $\tau_{\mathrm{anc}}=1$ → standard cross-entropy
    • KD: $q=\mathrm{softmax}(z_{T}/T)$ from teacher logits, $\tau_{\mathrm{anc}}=T$ → temperature-scaled KD
  • Fixed anchor: Set $\pi_{\mathrm{ref}}$ once (e.g., pretrained model, SFT checkpoint) and never update. The coordinate system remains fixed, providing a stable geometric center for:
    • Aligning to labels/rewards/preferences
    • Regularizing toward a known good policy
    • Preventing catastrophic forgetting
Key property: With fixed/no anchor, the optimization has monotonic convergence guarantees (Theorem 3.10).
Online ADPO: Dynamic anchor.
In online settings (RL, iterative RLHF, continual learning), the anchor is dynamically updated to track the evolving policy:
$\boxed{\mathcal{L}_{\text{ADPO-online}}\;=\;\mathbb{E}_{x}\Big[\mathrm{KL}\!\Big(q(\cdot|x)\ \Big\|\ \mathrm{softmax}\!\big((\ell_{\theta}-\ell_{\text{old}})/\tau_{\mathrm{anc}}\big)\Big)\Big],\quad\pi_{\mathrm{ref}}\leftarrow\pi_{\text{old}}.}$ (17)
(The anchored logits $(\ell_{\theta}-\ell_{\text{old}})/\tau_{\mathrm{anc}}$ are $q$-centered as in Eq. 10.)
Anchor update strategies:
  • Per-iteration: $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$ every iteration (like TRPO [23]/PPO [4])
  • Periodic: $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$ every $N$ iterations (balances stability/adaptivity)
  • EMA: $\pi_{\mathrm{ref}}\leftarrow\alpha\pi_{\mathrm{ref}}+(1-\alpha)\pi_{\theta}$ (smooth evolution)